The following data represents the weight in pounds of a sample of young boys that are between the ages of 13 years and 15 years

126    231    132    126   128   128    131   136    140    226    156    149 115 117    138    133   131   134    

  1. Make a frequency table that represents this data
  2. Calculate the mean an the median
  3. Calculate standard deviation
  4. Create intervals
  5. Interpret the result
Answer #1

ANSWER:

The following data represents the weight in pounds of a sample of young boys that are between the ages of 13 years and 15 years.

115,117,126,126,128,128,131,132,133,134,136,138,140,149,156,226,231.

Data Frequency
115 1
117 1
126 2
128 1
131 2
132 1
133 1
134 1
136 1
138 1
140 1
149 1
156 1
226 1
231 1

Mean=(sum of the items)/(no.of items)=2409/17=144.0588

Median=133

standard deviation=33.3119

Interval==[144.0588-33.3119,144.0588+33.3119]

Interval=[110.7469.177.3707].